The sum of two numbers is 20. The difference is 6
Answer by algebrahouse.com(1659)

x = one number
y = other number

x + y = 20 {sum of two numbers is 20}
x - y = 6 {their difference is 6}
----------
2x = 26 {added the two equations}
x = 13 {divided each side by 2}

x + y = 20 {first equation}
13 + y = 20 {substituted 13, in for x, into first equation}
y = 7 {subtracted 13 from each side}

13 and 7 are the two numbers

For more help from me, visit: www.algebrahouse.com